Cinnabar eyes (cn) and reduced bristles (rd) are autosomal recessive characters in Drosophila. A homozygous wild-type female was crossed to a reduced, cinnabar male, and the F1 males were then crossed to the F1 females to obtain the F2. Of the 400 F2 offspring obtained, 292 were wild type, 9 were cinnabar, 7 were reduced, and 92 were reduced and cinnabar.
Answer #1

Gene distance between cn and rd = number of recombinants / total number of progenies × 100

Number of recombinants = 9 + 7 = 16

Total number of progenies = 9 + 7 + 292 + 92 = 400

Gene distance (recombination frequency) = 16/400 × 100 = 4 mu

Please rate high.